• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 15:53:17 on November 11, 1998

    Entry number 5214

    keyword=Beam Time Accounting
    Hall A

    11/11/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: Suleiman
    Shift Workers: Chudakov, Incerti, Jutier, LeRose

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 15 1 1.25
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    1 1 1 1 1 1 45 0 6.75
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: - NMR at 10:25 : P = 32.6 (H2O) to 37 % (EPR) -08:00-13:20 CW Beam Loss Monitor (BLM) repairs -13:20 1uA CW on target. -13:40 15 uA. Start taking data Run # 2486 [parallel on Helium] -13:50 MCC kept getting BLM trips. They will go to pulsed beam -14:40 15 uA on target. Start data taking. Each 5 minutes the beam trips. -14:45 Run # 2488 is the first good run. -15:00 good beam.

    Data Entry by: Suleiman